# SoftShadow: Leveraging Penumbra-Aware Soft Masks for Shadow Removal

### v1

@@ 3.3 Penumbra Formation Constraint
Figure 5: Examples of shadow removal results on SRD datasets (Qu et al. 2017). The input shadow image, the estimated results of (a) DC-ShadowNet (Jin, Sharma, and Tan 2021), (b) DeS3 (Jin et al. 2024), and (c) Ours, as well as the ground truth image, respectively.
Figure 6: Examples of shadow removal results on the LRSS dataset (Gryka, Terry, and Brostow 2015). The input shadow image, the estimated results of (a) DC-ShadowNet (Jin, Sharma, and Tan 2021), (b) G2R-ShadowNet (Liu et al. 2021), (c) Ours, our predicted soft shadow masks and ground truth image, respectively.
Figure 7: Examples of soft shadow image removal results on the SRD dataset (Qu et al. 2017). The input shadow image, the estimated results of (a) BMNet (Zhu et al. 2022), (b) ShadowDiffusion (Guo et al. 2023b), (c) Homoformer (Xiao et al. 2024), and (d) Ours, as well as the ground truth image, respectively.
4 Experiments
4.1 Experimental Setups
Implementation details
The shadow removal backbone in our SoftShadow framework can be any shadow removal network. We use ShadowDiffusion (Guo et al. 2023b) as the backbone example for all experiments. We empirically set the threshold $t=0.76$ to get the ground truth soft shadow mask $\mathbf{s}_{gt}$. We employ the ViT-H model as the backbone for SAM (Kirillov et al. 2023). A set of LoRA (Hu et al. 2021) layers with a rank of 8 is added to the self-attention blocks in the image encoder of SAM. We set $\lambda_{1}=0.1$ and $\lambda_{2}=1$. For evaluation, we use the DDIM sampler (Song, Meng, and Ermon 2020) and 5 diffusion sampling steps. The training batch size is 16. For more training details, please refer to the supplementary.
Benchmark datasets We work with four benchmark datasets for the various shadow removal experiments. SRD Dataset (Qu et al. 2017) consists of 2,680 training pairs and 408 testing pairs of shadow and shadow-free images. Notably, the SRD dataset does not provide shadow masks, the previous methods using SRD commonly using masks detected by DHAN (Cun, Pun, and Shi 2020) methods. LRSS Dataset (Gryka, Terry, and Brostow 2015) is a specifically designed soft shadow dataset that includes 137 images. We select 48 paired shadow and shadow-free images as our testing set. UIUC Dataset (Guo, Dai, and Hoiem 2012) contains 76 pairs of images for testing. It features a variety of shadow types (Jin et al. 2024), including soft, hard, and self shadows, which provides a diverse and challenging set of conditions for evaluating shadow removal methods. ISTD+ Dataset (Le and Samaras 2019) is adjusted ISTD dataset (Wang, Li, and Yang 2018), reducing the illumination inconsistency between shadow and shadow-free images of the ISTD dataset. It manually annotated binary masks.
Evaluation metrics Following previous works (Cun, Pun, and Shi 2020; Guo et al. 2023b; Xiao et al. 2024), we employ the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) (Wang et al. 2004), and Mean Absolute Error (MAE) (Willmott and Matsuura 2005) as quantitative evaluation metrics. We also calculate all metrics for shadow areas, non-shadow areas, and all pixels between ground truth shadow-free images and generated removal results.
4.2 Comparison with State-of-the-Art
We compare our proposed method with several state-of-the-art shadow removal methods. We include methods that do not require shadow masks as input, e.g., DC-shadowNet (Jin, Sharma, and Tan 2021), DeS3 (Jin et al. 2024), and SAM-helps-shadow (Zhang, Gu, and Zhu 2023). And methods that require shadow masks, including DHAN (Cun, Pun, and Shi 2020), BMNet (Zhu et al. 2022), ShadowFormer (Guo et al. 2023a), Inpaint4Shadow (Li et al. 2023), ShadowDiffusion (Guo et al. 2023b) and Homoformer (Xiao et al. 2024). To demonstrate the generalizability of our method, we also compare it with zero-shot methods including BCDiff (Guo et al. 2023c) and G2R-ShadowNet (Liu et al. 2021) on the LRSS (Gryka, Terry, and Brostow 2015) and UIUC (Guo, Dai, and Hoiem 2012) datasets.
Quantitative results Table 1 shows the quantitative results of SRD and ISTD+ datasets. The SRD dataset contains many soft shadow images, which is more suitable for validating our methods. Specifically, we outperform all competing methods over all metrics, whether they require input shadow masks or not. Compared with the most recent work without requiring mask input, i.e., DeS3 (Jin et al. 2024), the PSNR is improved from 34.11 dB to 35.57 dB in whole images. When comparing with methods that need input masks, we outperform the SOTA method Homoformer (Xiao et al. 2024), which uses DHAN (Cun, Pun, and Shi 2020) mask as their input. On the ISTD+ dataset, we significantly outperform DeS3 (Jin et al. 2024), increasing the PSNR from 31.38 dB to 35.00 dB. For methods that rely on the ground truth masks from the ISTD+ dataset, we evaluate them using masks detected by the FDRNet (Zhu et al. 2021) followed by previous methods BMNet (Zhu et al. 2022), to ensure a fair comparison. We outperform all competing methods under the condition of without ground truth manually annotated masks. Besides, we even achieve comparable results against some SOTA methods using ground truth masks as shown in Table 1. To further evaluate the boundary artifacts suppress ability on SRD dataset, we present the restoration results in the penumbra area in Table 3. The penumbra area is obtained using the method from AEF (Fu et al. 2021). For methods that require binary masks as input, such as Homoformer, our penumbra area provides much better results.
@@ 4.2 Comparison with State-of-the-Art
Table 2: The quantitative results of shadow removal using our SoftShadow and recent methods on the LRSS dataset. Here, “DC” refer to DC-ShadowNet (Jin, Sharma, and Tan 2021), “G2R” refer to G2R-ShadowNet. The best results are boldfaced.
Methods PSNR MAE
@@
Table 3: The quantitative results in the penumbra area. The best results are boldfaced.
@@
SAM 21.78 0.913 12.07 35.27 0.973 3.32
$\mathcal{L}_{mask}$ 23.08 0.935 9.97 35.44 0.974 3.17
Ours 23.32 0.933 9.77 35.57 0.975 3.11
Table 4: The ablation studies for the mask reconstruction loss $\mathcal{L}_{mask}$ and the constraint loss $\mathcal{L}_{pen}$ in LRSS and SRD datasets. Here, the “SAM” means pretrained SAM. The best results are boldfaced.
Table 2 shows the generalizability of our methods. We use our pretrained model on the SRD dataset and test it on LRSS and UIUC datasets without further training. The DeS3 (Jin et al. 2024) has better results because they train their method on the LRSS training set. For comparison, we chose DC-ShadowNet as our baseline since this method does not require masks as input. Our results significantly outperform DC-ShadowNet on LRSS and UIUC datasets. Additionally, we compare our methods with two zero-shot methods, which are designed to have better generalizability. The results demonstrate that our method has the best PSNR/SSIM/MAE among the comparison methods.
Qualitative evaluation To further demonstrate the effectiveness of our methods compared to other competing methods, Figure 7 illustrates our improvements in removing soft shadows on the SRD dataset compared with methods using binary shadow masks. The soft masks typically appear as soft boundaries of large object shadows or in small shadow regions. Specifically, our samples demonstrate that we effectively remove both types of soft shadows while preserving the illumination transition in penumbra areas. Figure 5 shows our improvement in shadow area accuracy compared to other end-to-end methods. Our approach effectively removes all parts of the shadow shown in the first row. In the second row, we successfully remove shadows in regions with complex textures. Figure 6 highlights our method’s generalizability on the LRSS dataset. The results demonstrate that our method achieves better shadow removal performance without further training on the LRSS dataset.
Figure 8: The quantitative results of shadow removal and their predicted masks with Pretrained SAM, $\mathcal{L}_{mask}$, and $\mathcal{L}_{mask}+\mathcal{L}_{pen}$ (Ours), respectively.
5 Ablation Study
We ablate our method, SoftShadow, by removing two losses: the mask reconstruction loss $\mathcal{L}_{mask}$ and the penumbra gradient loss $\mathcal{L}_{pen}$, and demonstrate the results on SRD and LRSS datasets. First, we finetune the framework using only the shadow removal loss $\mathcal{L}_{rem}$. The results show that without shadow mask guidance, SAM struggles to detect shadow masks in some cases. Next, we add the mask reconstruction loss $\mathcal{L}_{mask}$ to provide SAM with more precise mask position guidance. The results indicate that the $\mathcal{L}_{mask}$ improves shadow removal performance. Finally, we added the penumbra gradient loss $\mathcal{L}_{pen}$ to refine the mask boundary. According to Table 4, we can see that in the LRSS dataset, both $\mathcal{L}_{mask}$ and $\mathcal{L}_{pen}$ significantly improve performance, but in the SRD dataset, $\mathcal{L}_{pen}$ does not show much improvement in terms of PSNR. We hypothesize that it might be caused by the proportion of soft areas in the SRD dataset is relatively small, so metrics like PSNR, which measure the average error, may not reflect the improvements. As shown in Figure 8, adding the constraint loss $\mathcal{L}_{pen}$ results in more accurate and smoother soft mask contour compared to using only $\mathcal{L}_{mask}$, leading to better results in soft shadow areas. From the perspective of the removal network, the constraint loss provides more detailed soft boundary information, leading to better results in soft shadow areas.

### v2

@@ 3.2 Architecture of SoftShadow
Moreover, as illustrate in Figure 3, we incorpotare the gradient of the shadow removal loss $\mathcal{L}_{rem}$ to jointly train SAM and the subsequent shadow removal network:
$\mathcal{L}_{rem}=\mathbb{E}_{n}\left\|\mathbf{\hat{x}}-\mathbf{x}\right\|_{F}^{2},$ (4)
where $n$ indexes all images, $\hat{\mathbf{x}}$ is the restored shadow-free image, and $\mathbf{x}$ is the ground truth shadow-free image. To this end, the intermediate soft shadow masks can be refined using shadow removal results, with the SAM adapter optimized for better location guidance in subsequent shadow removal.
The overall training objective of SoftShadow is:
where $\lambda_{1}$ and $\lambda_{2}$ are weighting coefficients to balance the influence of each term. With the merits of joint training and physical constraints, our SoftShadow framework effectively removes shadows.
Figure 4: Illustrate the concept of Penumbra Formation Constraint. The $w_{0}$ and $w_{1}$ are points in the inner boundary and the outer boundary of the penumbra area, respectively. In the ideal soft shadow mask, the intensity in the penumbra area decreases progressively from the shadow center to the umbra area. The gradient orientation points from the shadow center towards the non-shadow area.
3.3 Penumbra Formation Constraint
We introduce the Penumbra Formation Constraint to enable SAM to predict the soft masks more accurately. The objective of this constraint is to regularize SAM to learn a mask that can reflect the illuminance change of shadow images. To meet the objective, we present two assumptions for our proposed soft mask. For a given shadow image, the soft mask intensity in the penumbra region should gradually transition outward from the shadow center to the outer edge of the shadow. This means that the gradient direction in the penumbra should point from the shadow center toward the boundary of the shadow. Additionally, since the intensity change in this area is gradual, the gradient magnitude should remain relatively small to reflect a smooth transition. Based on these assumptions, we can formulate the penumbra formation constraint. For a predicted soft mask $\mathbf{s}$, we define the penumbra area using two thresholds $t_{1}$ and $t_{2}$ as
$\mathbf{w}=\{(i,j)\mid t_{1}\leq\mathbf{s}_{i,j}\leq t_{2}\},$ (6)
where $i$ and $j$ denote the spatial coordinates of pixels within the soft mask $\mathbf{s}$. As shown in Figure 4, the center of the shadow area is denoted as $\mathbf{c}$. $\mathbf{c}$ is defined as the mean of the $x$- and $y$- coordinates of the penumbra area. The unit vector representing the direction from $\mathbf{c}$ to $\mathbf{w}$ in the penumbra area can be expressed as:
$\mathbf{d}(\mathbf{w})=\frac{\mathbf{w}-\mathbf{c}}{\|\mathbf{w}-\mathbf{c}\|},\quad\text{for }\mathbf{w}\neq\mathbf{c}$ (7)
Following the assumptions we proposed, we can define the penumbra formation constraint loss $\mathcal{L}_{pen}$ as:
$\mathcal{L}_{\text{pen}}=\mathbb{E}_{n,\mathbf{w}}\left[R\left(\mathbf{d}(\mathbf{w})\cdot\nabla M(\mathbf{w})\right)\right],$ (8)
where $n$ is the index of shadow masks, the desired direction of gradient is $\mathbf{d}(\mathbf{w})$, and $\cdot$ indicate the dot product. $R$ refers to the ReLU function [25]. The ReLU function is applied to filter out gradients conflict with the desired direction. The constraint loss regularizes the magnitude of gradients in alignment with the desired direction, promoting a smooth intensity transition within the penumbra region. By penalizing quick changes in intensity, this constraint encourages the detector to generate softer shadow masks.
Methods Input Masks shadow non-shadow all
[…]
SRD DHAN DHAN 33.67 0.978 8.94 34.79 0.979 4.80 30.51 0.949 5.67
DC-ShadowNet N/A 34.00 0.975 7.70 35.53 0.981 3.65 31.53 0.955 4.65
BMNet DHAN 35.05 0.981 6.61 36.02 0.982 3.61 31.69 0.956 4.46
SG-ShadowNet DHAN 33.73 0.979 7.53 36.18 0.982 2.97 31.16 0.952 4.23
ShadowFormer DHAN 35.55 0.982 6.14 36.82 0.983 3.54 32.46 0.957 4.28
DeS3 N/A 37.91 0.986 5.27 37.45 0.984 3.03 34.11 0.968 3.56
Homoformer DHAN 38.81 0.987 4.25 39.45 0.988 2.85 35.37 0.972 3.33
SAM-helps-shadow N/A 33.94 0.979 7.44 33.85 0.981 3.74 30.72 0.952 4.79
ours N/A 39.08 0.989 4.33 39.36 0.992 2.58 35.57 0.975 3.11
ISTD+ BMNet GT 37.87 0.991 5.62 37.51 0.985 2.45 33.98 0.972 2.97
ShadowDiffusion GT 39.69 0.992 4.97 38.89 0.987 2.28 35.67 0.975 2.72
HomoFormer GT 39.49 0.993 4.73 38.75 0.984 2.23 35.35 0.975 2.64
DC-ShadowNet N/A 31.06 0.976 12.62 27.03 0.961 6.82 25.03 0.926 7.77
DeS3 N/A 36.49 0.989 6.56 34.70 0.972 3.40 31.38 0.958 3.94
BMNet FDRNet - - 6.1 - - 2.9 - - 3.5
HomoFormer FDRNet 38.84 0.991 5.31 34.58 0.966 3.17 32.41 0.953 3.51
Ours N/A 40.36 0.993 4.78 37.89 0.982 2.46 35.00 0.972 2.85
Table 1: The quantitative results of shadow removal using our SoftShadow and recent methods on SRD and ISTD+ datasets. The “Input Masks” column shows the different types of input masks used by these methods. “N/A” means the method does not require masks as input. “GT” means the method uses manually annotated ground truth mask as input. “DHAN” means using masks generated by DHAN [3] method. “FDRNet” means using masks generated by FDRNet [39] method. The best and the second results are boldfaced and underlined, respectively.
@@ 3.3 Penumbra Formation Constraint
Table 2: The quantitative results of shadow removal using our SoftShadow and recent methods on the LRSS and UIUC datasets. The best results are boldfaced.
[…]
@@
Table 3: The quantitative results in the penumbra area. The penumbra area is calculated from our predicted soft masks. The best results are boldfaced.
Input Masks ShadowDiffusion HomoFormer Ours
Ground Truth mask 35.67 35.35 35.00
Otsu mask 35.62 33.47
BDRAR mask 33.41 32.96
FDRNet mask 34.08 32.41
Pretrained SAM mask 28.45 28.16
Mean 33.45 32.47 N/A
Std Dev 0.912 0.993 N/A
Table 4: The mask sensitivity evaluation on ISTD+ dataset. The values in the table are PSNR results. The best results are boldfaced. The results that are higher than ours are underlined.
4 Experiments
4.1 Experimental Setups
Implementation details
The shadow removal backbone in our SoftShadow framework can be any shadow removal network. We use ShadowDiffusion [9] as the backbone example for all experiments. We finetune ShadowDiffusion at a resolution of 256 $\times$ 256, followed by previous methods [9]. We empirically set the threshold $t=0.76$ to get the ground truth soft shadow mask $\mathbf{s}_{gt}$. We employ the ViT-H model as the backbone for SAM [19]. A set of LoRA [13] layers with a rank of 8 is added to the self-attention blocks in the image encoder of SAM. We use Adam [18] optimizer. The training batch size is 16. We set $\lambda_{1}=0.1$ and $\lambda_{2}=1$. For evaluation, we use the DDIM sampler [29] and 5 diffusion sampling steps. For more training details, please refer to the supplementary.
Benchmark datasets We work with four benchmark datasets for the various shadow removal experiments. SRD Dataset [27] consists of 2,680 training pairs and 408 testing pairs of shadow and shadow-free images. Notably, the SRD dataset does not provide shadow masks, the previous methods using SRD commonly using masks detected by DHAN [3] methods. LRSS Dataset [7] is a specifically designed soft shadow dataset that includes 137 images. We select 48 paired shadow and shadow-free images as our testing set. UIUC Dataset [11] contains 76 pairs of images for testing. It features a variety of shadow types [17], including soft, hard, and self shadows, which provides a diverse and challenging set of conditions for evaluating shadow removal methods. ISTD+ Dataset [21] is adjusted ISTD dataset [32], consists of 1330 training pairs and 540 testing pairs of shadow and shadow-free images. ISTD+ has manually annotated binary masks.
Evaluation metrics Following previous works [3, 9, 35], we employ the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [33], and Mean Absolute Error (MAE) [34] as quantitative evaluation metrics. We also calculate all metrics for shadow areas, non-shadow areas, and all pixels between ground truth shadow-free images and generated removal results.
4.2 Comparison with State-of-the-Art
We compare our proposed method with several state-of-the-art shadow removal methods. We include methods that do not require shadow masks as input, e.g., DC-shadowNet [16], DeS3 [17], and SAM-helps-shadow [36]. And methods that require shadow masks, including DHAN [3], BMNet [40], SG-ShadowNet [31], ShadowFormer [8], Inpaint4Shadow [22], ShadowDiffusion [9] and Homoformer [35]. To demonstrate the generalizability of our method, we also compare it with zero-shot methods including BCDiff [10] and G2R-ShadowNet [24] on the LRSS [7] and UIUC [11] datasets.
Quantitative results Table 1 shows the quantitative results of SRD and ISTD+ datasets. The SRD dataset contains many soft shadow images, which is more suitable for validating our methods. Specifically, we outperform all competing methods over all metrics, whether they require input shadow masks or not. Compared with the most recent work without requiring mask input, i.e., DeS3 [17], the PSNR is improved from 34.11 dB to 35.57 dB in whole images. When compared with methods that need input masks, we outperform the SOTA method Homoformer [35], which uses DHAN [3] mask as their input. On the ISTD+ dataset, we significantly outperform DeS3 [17], increasing the PSNR from 31.38 dB to 35.00 dB. For methods that rely on the ground truth masks from the ISTD+ dataset, we evaluate them using masks detected by the FDRNet [39] followed by previous methods BMNet [40], to ensure a fair comparison. We outperform all competing methods under the condition of without ground truth manually annotated masks. Besides, we even achieve comparable results against some SOTA methods using ground truth masks.
Table 2 shows the generalizability of our methods. We use our pretrained model on the SRD dataset and test it on LRSS and UIUC datasets without further training. The DeS3 [17] has better results because they train their method on the LRSS training set. For comparison, we chose DC-ShadowNet as our baseline since this method does not require masks as input. Our results significantly outperform DC-ShadowNet on LRSS and UIUC datasets. Additionally, we compare our methods with two methods, which are designed to have better generalizability. The results demonstrate that our method has the best PSNR/SSIM/MAE among the comparison methods. These results highlight the robustness and adaptability of our model, even on datasets it was not specifically trained on.
Figure 5: Examples of soft shadow image removal results on the SRD dataset [27]. The input shadow image, the estimated results of (a) BMNet [40], (b) ShadowDiffusion [9], (c) Inpaint4Shadow [22], (d) Homoformer [35], and (e) Ours, as well as the ground truth (GT) image, respectively.
Figure 6: Examples of soft shadow removal results on SRD datasets [27]. The input shadow image, the estimated results of (a) DC-ShadowNet [16], (b) DeS3 [17], and (c) Ours, as well as the ground truth image, respectively.
Figure 7: Examples of shadow removal results on the LRSS dataset [7]. The input shadow image, the estimated results of (a) DC-ShadowNet [16], (b) G2R-ShadowNet [24], (c) Ours, our predicted soft shadow masks and ground truth image, respectively.
Qualitative evaluation To further demonstrate the effectiveness of our methods compared to other competing methods, Figure 5 illustrates our improvements in removing soft shadows on the SRD dataset compared with methods using hard shadow masks. The soft masks typically appear as soft boundaries of large object shadows or in small shadow regions. Specifically, our samples demonstrate that we effectively remove both types of soft shadows while preserving the illumination transition in penumbra areas, without introducing obvious artifacts. Figure 6 shows our improvement in shadow area accuracy compared to other end-to-end methods. Our approach effectively removes all parts of the shadow shown in the first row of the example. In the second and third rows, we successfully remove shadows in regions with complex textures. Figure 7 highlights our method’s generalizability on the LRSS dataset, achieving improved shadow removal performance without additional training on LRSS dataset. These results confirm our method’s adaptability and precision in diverse shadowed scenarios.
5 Framework Analysis
In this section, we provide more experiments to evaluate the effectiveness of our proposed SoftShadow framework. We discuss the performance on the penumbra area, the mask sensitive of previous methods, and we discuss the effectiveness of our provided losses in ablation study.
Penumbra effect evaluation The penumbra region typically occupies only a small portion of the entire image, so improvements within this area have a limited impact on overall image metrics. To further assess the improvements introduced by using soft masks in the penumbra region, we calculate the PSNR and MAE values specifically within this area on the SRD dataset. For a fair comparison, we define the penumbra region following the methods from AEF [5], applying soft masks generated by our method. The results are shown in Table 3. Compared with the SOTA methods like HomoFormer [35], DeS3 [17] and Inpaint4Shadow [22], we have improvements on PSNR for around 1db. These results show that on images containing soft shadows, our methods has better restoration performances.
Mask sensitivity evaluation Previous shadow removal methods that rely on external shadow masks are sensitive to the accuracy of the input masks. Their performance can degrade when lack of ground truth mask. In contrast, end-to-end shadow removal methods mitigate the impact of low-quality masks. To further illustrate the sensitivity of prior SOTA methods to input masks, we evaluate their performance using various mask inputs.
Specifically, we evaluate ShadowDiffusion [9] and HomoFormer [35] using a range of input masks, including ground truth (GT) masks, Otsu masks [14], BDRAR [38] detected masks, FDRNet [39] detected masks and pretrained SAM [19] detected masks as inputs. The PSNR results on the ISTD+ [21] datasets are shown in Table 4. The Otsu mask follows the approach in Mask-ShadowGAN [14], using the difference between shadow-free and shadowed images to generate masks; both Otsu and GT masks include additional shadow-free information. Results indicate that our methods maintain consistent performance across all tests, whereas ShadowDiffusion and HomoFormer have performance declines when using generated masks. Although our method achieves slightly lower PSNR values than those obtained with Otsu or GT masks, this is reasonable given that the ISTD+ dataset contains a larger proportion of hard shadow images, where our method has less of an advantage. Our approach surpasses the rest results that rely on masks generated by shadow detectors.
Ablation Study We ablate our method, SoftShadow, by adding three losses: the shadow removal loss $\mathcal{L}_{rem}$, the mask reconstruction loss $\mathcal{L}_{mask}$, and the penumbra formation constraint loss $\mathcal{L}_{pen}$, and demonstrate the results on SRD and LRSS datasets. First, we present the shadow removal results using the pretrained SAM and ShadowDiffusion models. Next, we finetune the framework with the shadow removal loss $\mathcal{L}_{rem}$. Then, we add the mask reconstruction loss $\mathcal{L}_{mask}$ to provide SAM with more precise mask position guidance. Finally, we add the penumbra formation constraint loss $\mathcal{L}_{pen}$ to refine the mask boundaries for smoother transitions in the shadow regions.
The results in Table 5 are based on training with the SRD training set and evaluating on both the SRD and LRSS datasets. As shown in Table 5, the shadow removal loss $\mathcal{L}_{rem}$ improves performance on the SRD dataset, primarily by enhancing the shadow removal network. Adding the mask reconstruction loss $\mathcal{L}_{mask}$ further fine-tunes SAM, producing more accurate shadow masks. This improvement of mask accuracy leads to better shadow removal results. To evaluate the impact of the penumbra formation constraint loss $\mathcal{L}_{pen}$, we compare results with and without $\mathcal{L}_{pen}$. The results show that $\mathcal{L}_{pen}$ further aids in shadow removal on the SRD dataset, refining the shadow edges and enhancing the overall shadow removal quality. In the LRSS dataset, both $\mathcal{L}_{rem}$ and $\mathcal{L}_{mask}$ significantly improve performance, the $\mathcal{L}_{pen}$ improves the PSNR and MAE significantly, although it results in a slight decrease in SSIM. As shown in Figure 8, adding the constraint loss $\mathcal{L}_{pen}$ leads to more accurate and smoother soft mask predictions. The combined effects of $\mathcal{L}_{mask}$ and $\mathcal{L}_{pen}$ lead to more accurate and smoother soft mask predictions. Together, these losses provide the removal framework with enhanced soft shadow details, resulting in superior performance in shadow removal.
@@ 5 Framework Analysis
Pretrained Weights 21.40 0.910 12.26 31.27 0.963 4.41
$\mathcal{L}_{rem}$ 21.78 0.913 12.07 35.27 0.973 3.32
$\mathcal{L}_{rem}+\mathcal{L}_{mask}$ 23.08 0.935 9.97 35.44 0.974 3.17
$\mathcal{L}_{rem}+\mathcal{L}_{mask}+\mathcal{L}_{pen}$ 23.32 0.933 9.77 35.57 0.975 3.11
Table 5: The ablation studies for the shadow removal loss $\mathcal{L}_{rem}$, the mask reconstruction loss $\mathcal{L}_{mask}$, and the penumbra formation constraint loss $\mathcal{L}_{pen}$ in LRSS and SRD datasets. The best results are boldfaced.
Figure 8: Visual comparison of results from our framework with different configurations: using pretrained weights, finetuned with the loss function $\mathcal{L}_{rem}$, $\mathcal{L}_{rem}+\mathcal{L}_{mask}$ and our full configuration $\mathcal{L}_{rem}+\mathcal{L}_{mask}+\mathcal{L}_{pen}$ (Ours), respectively.